# Maximum Likelihood Estimation Based Complex-Valued Robust Chinese Remainder Theorem and Its Fast Algorithm

### v1

@@ III-B MLE C-CRT
Now, we calculate the maximum likelihood function for $\mathsf{N}\in\mathcal{F}_{M\Gamma}$, which is the fixed but unknown complex number to determine as mentioned above. Assume that $\mathsf{X}_{i}=\mathsf{N}+\mathsf{W}_{i}$ follow complex Gaussian distributions, and $\mathsf{W}_{i}$ are independent of each other, where $i=1,2,\ldots,L$. The variances of the real and imaginary parts of $\mathsf{W}_{i}$ are both $\sigma_{i}^{2}$, and their means are both 00 for each $i$. Denote $\mathsf{R}_{i}$ as the random variable with observed value $\tilde{\mathsf{r}}_{i}$, which satisfies $\mathsf{R}_{i}\equiv\mathsf{X}_{i}\mod M\mathsf{\Gamma}_{i}$ and $\Gamma=\prod_{i=1}^{L}\mathsf{\Gamma}_{i}$ is an integer as assumed earlier. Then, $\mathsf{R}_{i}$ follows a wrapped complex Gaussian distribution as studied above. For each $i$, from (19), we can obtain the conditional pdf of $\mathsf{R}_{i}$ as follows:
@@
Generally, $|M\mathsf{\Gamma}_{i}|$ is much larger than $\sigma_{i}$. By Proposition 2 and (24), we have $f_{\mathsf{R}_{i}}(\tilde{\mathsf{r}}_{i}\mid\mathsf{N})\approx\frac{1}{2\pi\sigma_{i}^{2}}\exp\left\{-\frac{|d_{M\mathsf{\Gamma}_{i}}(\tilde{\mathsf{r}}_{i},\mathsf{N})|^{2}}{2\sigma_{i}^{2}}\right\}$. Consequently, we can approximate the joint conditional pdf $f_{\mathsf{R}_{1},\mathsf{R}_{2},\ldots,\mathsf{R}_{L}}(\tilde{\mathsf{r}}_{1},\tilde{\mathsf{r}}_{2},\ldots,\tilde{\mathsf{r}}_{L}\mid\mathsf{N})=\prod_{i=1}^{L}f_{\mathsf{R}_{i}}(\tilde{\mathsf{r}}_{i}\mid\mathsf{N})$ as
@@
$\textstyle\hat{\mathsf{N}}_{\text{MLE}}$ $=\textstyle\arg\max\limits_{\mathsf{z}\in\mathcal{F}_{M\Gamma}}\mathcal{L}\left(\mathsf{z}\right)$ (26)
$=\textstyle\arg\min\limits_{\mathsf{z}\in\mathcal{F}_{M\Gamma}}\sum_{i=1}^{L}{\frac{1}{\sigma_{i}^{2}}|d_{M\mathsf{\Gamma}_{i}}\left(\tilde{\mathsf{r}}_{i},\mathsf{z}\right)|^{2}}.$
@@ III-C Fast MLE C-CRT Algorithm
By (29) and Theorem 1, we have $\hat{\mathsf{N}}_{0}\equiv\hat{\mathsf{q}}_{i}\mod\mathsf{\Gamma}_{i}$ for $i=1,2,\ldots,L$. That is, there exist Gaussian integers $\mathsf{k}_{i}$ such that $\hat{\mathsf{N}}_{0}=\mathsf{k}_{i}\mathsf{\Gamma}_{i}+\hat{\mathsf{q}}_{i}$. By (30), we have $\hat{\mathsf{N}}=M\mathsf{k}_{i}\mathsf{\Gamma}_{i}+M\hat{\mathsf{q}}_{i}+\hat{\mathsf{r}}^{c}$. According to Property 4, we have $|d_{M\mathsf{\Gamma}_{i}}(\tilde{\mathsf{r}}_{i},\hat{\mathsf{N}})|=\left|d_{M\mathsf{\Gamma}_{i}}(\tilde{\mathsf{r}}_{i},M\hat{\mathsf{q}}_{i}+\hat{\mathsf{r}}^{c})\right|=\left|d_{M\mathsf{\Gamma}_{i}}(\tilde{\mathsf{r}}_{i}-M\hat{\mathsf{q}}_{i}-\hat{\mathsf{r}}^{c},0)\right|$. It follows from (28) that $|d_{M\mathsf{\Gamma}_{i}}(\tilde{\mathsf{r}}_{i},\hat{\mathsf{N}})|=\left|d_{M\mathsf{\Gamma}_{i}}(d_{M}(\tilde{\mathsf{r}}_{i},\hat{\mathsf{r}}^{c}),0)\right|$. As $d_{M}\left(\tilde{\mathsf{r}}_{i},\hat{\mathsf{r}}^{c}\right)\in\mathcal{S}_{M}$, by Property 7, we have $|d_{M\mathsf{\Gamma}_{i}}(\tilde{\mathsf{r}}_{i},\hat{\mathsf{N}})|=\left|d_{M}(\tilde{\mathsf{r}}_{i},\hat{\mathsf{r}}^{c})\right|$. Hence, (26) and (27) are equivalent, that is, $\mathsf{z}$ in (26) is optimal if and only if $\mathsf{x}$ in (27) is optimal. ∎
@@
$w_{i}=\frac{\sigma_{i}^{-2}}{\sum_{i=1}^{L}\sigma_{i}^{-2}},\ i=1,2,\ldots,L.$ (31)
@@
$&\textstyle\mathrm{Re}(\hat{\mathsf{r}}^{c})=\arg\min\limits_{\mathrm{Re}(\mathsf{x})\in[0,M)}\sum_{i=1}^{L}w_{i}d_{M}^{2}\left(\mathrm{Re}(\tilde{\mathsf{r}}_{i}^{c}),\mathrm{Re}(\mathsf{x})\right),\\ &\textstyle\mathrm{Im}(\hat{\mathsf{r}}^{c})=\arg\min\limits_{\mathrm{Im}(\mathsf{x})\in[0,M)}\sum_{i=1}^{L}w_{i}d_{M}^{2}\left(\mathrm{Im}(\tilde{\mathsf{r}}_{i}^{c}),\mathrm{Im}(\mathsf{x})\right).\end{aligned}\right.$ (33)
@@
$\textstyle\sum_{i=1}^{L}w_{i}|d_{M}\left(\tilde{\mathsf{r}}_{i}^{c},\mathsf{x}\right)|^{2}=\sum_{i=1}^{L}w_{i}\left|\tilde{\mathsf{r}}_{i}^{c}-\mathsf{x}-\left[\frac{\tilde{\mathsf{r}}_{i}^{c}-\mathsf{x}}{M}\right]\right|^{2}$
$\textstyle=\sum_{i=1}^{L}w_{i}\left(\left(\mathrm{Re}(\tilde{\mathsf{r}}_{i}^{c})-\mathrm{Re}(\mathsf{x})-\left[\frac{\mathrm{Re}(\tilde{\mathsf{r}}_{i}^{c})-\mathrm{Re}(\mathsf{x})}{M}\right]\right)^{2}\right.$
$\textstyle\hskip 62.59596pt\left.+\left(\mathrm{Im}(\tilde{\mathsf{r}}_{i}^{c})-\mathrm{Im}(\mathsf{x})-\left[\frac{\mathrm{Im}(\tilde{\mathsf{r}}_{i}^{c})-\mathrm{Im}(\mathsf{x})}{M}\right]\right)^{2}\right)$
$\textstyle=\sum_{i=1}^{L}w_{i}d_{M}^{2}\left(\mathrm{Re}(\tilde{\mathsf{r}}_{i}^{c}),\mathrm{Re}(\mathsf{x})\right)+\sum_{i=1}^{L}w_{i}d_{M}^{2}\left(\mathrm{Im}(\tilde{\mathsf{r}}_{i}^{c}),\mathrm{Im}(\mathsf{x})\right).$
@@
[…]
$\textstyle k_{1},k_{2}=1,2,\ldots,L\Big{\}},$ (34)
where $\varsigma$ and $\upsilon$ are permutations on $\{1,2,\ldots,L\}$ satisfying $\mathrm{Re}(\tilde{\mathsf{r}}_{\varsigma_{(1)}}^{c})\leq\mathrm{Re}(\tilde{\mathsf{r}}_{\varsigma_{(2)}}^{c})\leq\cdots\leq\mathrm{Re}(\tilde{\mathsf{r}}_{\varsigma_{(L)}}^{c})$ and $\mathrm{Im}(\tilde{\mathsf{r}}_{\upsilon_{(1)}}^{c})\leq\mathrm{Im}(\tilde{\mathsf{r}}_{\upsilon_{(2)}}^{c})\leq\cdots\leq\mathrm{Im}(\tilde{\mathsf{r}}_{\upsilon_{(L)}}^{c})$, respectively.
@@ IV-A Condition of Robust Estimation
We first consider a necessary condition of robust estimation for the MLE C-CRT. For convenience, we define the remainder error set as $\mathcal{U}=\left\{\Delta\mathsf{r}_{1},\Delta\mathsf{r}_{2},\ldots,\Delta\mathsf{r}_{L}\right\}$ and the weighted average of the remainder errors as $\overline{\Delta\mathsf{r}}=\sum_{i=1}^{L}w_{i}\Delta\mathsf{r}_{i}$, where the weights $w_{i}$ are defined in (31). In [21], a necessary condition for a robust estimation of real numbers is $-\frac{M}{2}\leq\Delta r_{i}-(\hat{N}-N)<\frac{M}{2}$. For complex numbers, we have the following necessary condition:
@@
$\mathrm{Re}(\Delta\mathsf{r}^{c})=\mathrm{Re}(\overline{\Delta\mathsf{r}})+\begin{cases}M,&\text{if}~{}\mathrm{Re}(\mathsf{r}^{c}+\overline{\Delta\mathsf{r}})<0,\\ 0,&\text{if}~{}0\leq\mathrm{Re}(\mathsf{r}^{c}+\overline{\Delta\mathsf{r}})<M,\\ -M,&\text{if}~{}\mathrm{Re}(\mathsf{r}^{c}+\overline{\Delta\mathsf{r}})\geq M\end{cases}$ (40)
@@
$\mathrm{Im}(\Delta\mathsf{r}^{c})=\mathrm{Im}(\overline{\Delta\mathsf{r}})+\begin{cases}M,&\text{if}~{}\mathrm{Im}(\mathsf{r}^{c}+\overline{\Delta\mathsf{r}})<0,\\ 0,&\text{if}~{}0\leq\mathrm{Im}(\mathsf{r}^{c}+\overline{\Delta\mathsf{r}})<M,\\ -M,&\text{if}~{}\mathrm{Im}(\mathsf{r}^{c}+\overline{\Delta\mathsf{r}})\geq M.\end{cases}$ (41)
@@
For $\mathsf{N}$ and $\mathsf{N}_{0}$ in (14), $\mathsf{N}_{0}\in\mathcal{H}$ if and only if $\mathsf{N}\in M\mathcal{H}=\{M\mathsf{h}:\mathsf{h}\in\mathcal{H}\}$, where =$\mathcal{H}=\{h_{1}+h_{2}\mathrm{i}:1\leq h_{1},h_{2}<\Gamma-1\}$.
@@
Arbitrarily choose a $\Delta\mathsf{r}_{i}\in\mathcal{U}$, and let $\mathcal{V}=\{\Delta\mathsf{r}_{i}\}$. Then, we obtain from (38) that $\Delta\mathsf{r}_{i}-\sum_{j\neq i}\frac{w_{j}}{\sum_{j\neq i}w_{j}}\Delta\mathsf{r}_{j}\in\mathcal{S}_{M}$. Since $\sum_{j\neq i}w_{j}=1-w_{i}$, we have $\frac{1}{1-w_{i}}\Delta\mathsf{r}_{i}-\frac{1}{1-w_{i}}\overline{\Delta\mathsf{r}}\in\mathcal{S}_{M}$. Hence, $\Delta\mathsf{r}_{i}-\overline{\Delta\mathsf{r}}\in(1-w_{i})\mathcal{S}_{M}$. Consequently, $\left[\frac{\Delta\mathsf{r}_{i}-\overline{\Delta\mathsf{r}}}{M}\right]=0$. It follows from (42) that $\hat{\mathsf{q}}_{i}=\mathsf{q}_{i}-k_{1}-k_{2}\mathrm{i}$. Based on the arbitrariness of $\Delta\mathsf{r}_{i}$, we can obtain from (29) that
@@
Note that $\bar{\mathsf{\gamma}}_{i}\mathsf{\gamma}_{i}\equiv 1\mod\mathsf{\Gamma}_{i}$ and $\bar{\mathsf{\gamma}}_{j}\mathsf{\gamma}_{j}\equiv 0\mod\mathsf{\Gamma}_{i}$ for $j\neq i$. Hence, $\sum_{i=1}^{L}\bar{\mathsf{\gamma}}_{i}\mathsf{\gamma}_{i}\equiv 1\mod\mathsf{\Gamma}_{i}$. Consequently, $\sum_{i=1}^{L}\bar{\mathsf{\gamma}}_{i}\mathsf{\gamma}_{i}(k_{1}+k_{2}\mathrm{i})\equiv k_{1}+k_{2}\mathrm{i}\mod\mathsf{\Gamma}_{i}$, that is, $\mathsf{\Gamma}_{i}$ divides $\sum_{i=1}^{L}\bar{\mathsf{\gamma}}_{i}\mathsf{\gamma}_{i}(k_{1}+k_{2}\mathrm{i})-(k_{1}+k_{2}\mathrm{i})$. Since $\mathsf{\Gamma}_{1},\mathsf{\Gamma}_{2},\ldots,\mathsf{\Gamma}_{L}$ are pairwise coprime, $\prod_{i=1}^{L}\mathsf{\Gamma}_{i}$ divides $\sum_{i=1}^{L}\bar{\mathsf{\gamma}}_{i}\mathsf{\gamma}_{i}(k_{1}+k_{2}\mathrm{i})-(k_{1}+k_{2}\mathrm{i})$. Thus, $\sum_{i=1}^{L}\bar{\mathsf{\gamma}}_{i}\mathsf{\gamma}_{i}(k_{1}+k_{2}\mathrm{i})\equiv k_{1}+k_{2}\mathrm{i}\mod\Gamma$. By (43), we have $\hat{\mathsf{N}}_{0}\equiv\mathsf{N}_{0}-k_{1}-k_{2}\mathrm{i}\mod\Gamma$. Since $\hat{\mathsf{N}}_{0}\in\mathcal{F}_{\Gamma}$, we have $\hat{\mathsf{N}}_{0}=\left\langle\mathsf{N}_{0}-k_{1}-k_{2}\mathrm{i}\right\rangle_{\Gamma}$. If $\mathsf{N}\in M\mathcal{H}$, i.e., $\mathsf{N}_{0}\in\mathcal{H}$ by Proposition 3, then $\hat{\mathsf{N}}_{0}=\mathsf{N}_{0}-k_{1}-k_{2}\mathrm{i}$. Consequently, $\hat{\mathsf{N}}=M(\mathsf{N}_{0}-k_{1}-k_{2}\mathrm{i})+\mathsf{r}^{c}+\Delta\mathsf{r}^{c}=\mathsf{N}+\overline{\Delta\mathsf{r}}$. Therefore, $\mathsf{N}$ can be robustly estimated. The next theorem demonstrates that (39) is both a necessary and sufficient condition for robust estimation.
@@
Theorem 6 demonstrates that the MLE C-CRT is capable of “preserving errors”, i.e., it preserves the weighted average error $\overline{\Delta\mathsf{r}}$ of the original remainder errors $\Delta\mathsf{r}_{i},i=1,2,\ldots,L$, during the reconstruction. In this case, it is called error preserving MLE C-CRT. Since the output error of the MLE C-CRT is comparable to the input remainder error level, it is robust. By Theorem 6, we can derive the following robustness as well.
@@
Since $\mathrm{Re}(\Delta\mathsf{r}_{i})<\tau$ and $\mathrm{Im}(\Delta\mathsf{r}_{i})<\tau$, we have $\Big{|}\sum_{\Delta r_{i}\in S}\frac{w_{i}\mathrm{Re}(\Delta\mathsf{r}_{i})}{\sum_{\Delta r_{j}\in S}w_{j}}-\sum_{\Delta r_{i}\in\overline{S}}\frac{w_{i}\mathrm{Re}(\Delta\mathsf{r}_{i})}{\sum_{\Delta r_{j}\in\overline{S}}w_{j}}\Big{|}<\Big{|}\sum_{\Delta r_{i}\in S}\frac{w_{i}\tau}{\sum_{\Delta r_{j}\in S}w_{j}}\Big{|}+\Big{|}\sum_{\Delta r_{i}\in\overline{S}}\frac{w_{i}\tau}{\sum_{\Delta r_{j}\in\overline{S}}w_{j}}\Big{|}=2\tau$. Similarly, $\Big{|}\sum_{\Delta r_{i}\in S}\frac{w_{i}\mathrm{Im}(\Delta\mathsf{r}_{i})}{\sum_{\Delta r_{j}\in S}w_{j}}-\sum_{\Delta r_{i}\in\overline{S}}\frac{w_{i}\mathrm{Im}(\Delta\mathsf{r}_{i})}{\sum_{\Delta r_{j}\in\overline{S}}w_{j}}\Big{|}<2\tau$. By Theorem 6, we have $|\mathrm{Re}(\hat{\mathsf{N}})-\mathrm{Re}(\mathsf{N})|=|\mathrm{Re}(\overline{\Delta\mathsf{r}})|<\tau$ and $|\mathrm{Im}(\hat{\mathsf{N}})-\mathrm{Im}(\mathsf{N})|=|\mathrm{Im}(\overline{\Delta\mathsf{r}})|<\tau$. ∎

### v3

@@ III-B MLE C-CRT
Now, we calculate the maximum likelihood function for $\mathsf{N}\in\mathcal{F}_{M\Gamma}$, which is the fixed but unknown complex number to determine as mentioned above. Assume that $\mathsf{X}_{i}=\mathsf{N}+\mathsf{W}_{i}$ follow complex Gaussian distributions, and $\mathsf{W}_{i}$ are independent of each other, where $i=1,2,\ldots,L$. The variances of the real and imaginary parts of $\mathsf{W}_{i}$ are both $\sigma_{i}^{2}$, and their means are both 0 for each $i$. Denote $\mathsf{R}_{i}$ as the random variable with observed value $\tilde{\mathsf{r}}_{i}$, which satisfies $\mathsf{R}_{i}\equiv\mathsf{X}_{i}\mod M\mathsf{\Gamma}_{i}$ and $\Gamma=\prod_{i=1}^{L}\mathsf{\Gamma}_{i}$ is an integer as assumed earlier. Then, $\mathsf{R}_{i}$ follows a wrapped complex Gaussian distribution as studied above. For each $i$, from (20), we can obtain the conditional pdf of $\mathsf{R}_{i}$ as follows:
@@
Generally, $|M\mathsf{\Gamma}_{i}|$ is much larger than $\sigma_{i}$. By Proposition 2 and (25), we have
$f_{\mathsf{R}_{i}}(\tilde{\mathsf{r}}_{i}\mid\mathsf{N})\approx\frac{1}{2\pi\sigma_{i}^{2}}\exp\left\{-\frac{|d_{M\mathsf{\Gamma}_{i}}(\tilde{\mathsf{r}}_{i},\mathsf{N})|^{2}}{2\sigma_{i}^{2}}\right\}.$
Consequently, we can approximate the joint conditional pdf $f_{\mathsf{R}_{1},\mathsf{R}_{2},\ldots,\mathsf{R}_{L}}(\tilde{\mathsf{r}}_{1},\tilde{\mathsf{r}}_{2},\ldots,\tilde{\mathsf{r}}_{L}\mid\mathsf{N})=\prod_{i=1}^{L}f_{\mathsf{R}_{i}}(\tilde{\mathsf{r}}_{i}\mid\mathsf{N})$ as
@@
$\hat{\mathsf{N}}_{\text{MLE}}$ $=\arg\max\limits_{\mathsf{z}\in\mathcal{F}_{M\Gamma}}\mathcal{L}\left(\mathsf{z}\right)=\arg\min\limits_{\mathsf{z}\in\mathcal{F}_{M\Gamma}}\sum_{i=1}^{L}{\frac{1}{\sigma_{i}^{2}}|d_{M\mathsf{\Gamma}_{i}}\left(\tilde{\mathsf{r}}_{i},\mathsf{z}\right)|^{2}}.$ (27)
@@ III-C Fast MLE C-CRT Algorithm
By (30) and Theorem 1, we have
$\hat{\mathsf{N}}_{0}\equiv\hat{\mathsf{q}}_{i}\mod\mathsf{\Gamma}_{i},\ i=1,2,\ldots,L.$
That is, there exist Gaussian integers $\mathsf{k}_{i}$ such that $\hat{\mathsf{N}}_{0}=\mathsf{k}_{i}\mathsf{\Gamma}_{i}+\hat{\mathsf{q}}_{i}$. By (31), we have
$\hat{\mathsf{N}}=M\mathsf{k}_{i}\mathsf{\Gamma}_{i}+M\hat{\mathsf{q}}_{i}+\hat{\mathsf{r}}^{c}.$
According to Property 4, we have
$|d_{M\mathsf{\Gamma}_{i}}(\tilde{\mathsf{r}}_{i},\hat{\mathsf{N}})|=\left|d_{M\mathsf{\Gamma}_{i}}(\tilde{\mathsf{r}}_{i},M\hat{\mathsf{q}}_{i}+\hat{\mathsf{r}}^{c})\right|=\left|d_{M\mathsf{\Gamma}_{i}}(\tilde{\mathsf{r}}_{i}-M\hat{\mathsf{q}}_{i}-\hat{\mathsf{r}}^{c},0)\right|.$
It follows from (29) that
$|d_{M\mathsf{\Gamma}_{i}}(\tilde{\mathsf{r}}_{i},\hat{\mathsf{N}})|=\left|d_{M\mathsf{\Gamma}_{i}}(d_{M}(\tilde{\mathsf{r}}_{i},\hat{\mathsf{r}}^{c}),0)\right|.$
As $d_{M}\left(\tilde{\mathsf{r}}_{i},\hat{\mathsf{r}}^{c}\right)\in\mathcal{S}_{M}$, by Property 7, we have $|d_{M\mathsf{\Gamma}_{i}}(\tilde{\mathsf{r}}_{i},\hat{\mathsf{N}})|=\left|d_{M}(\tilde{\mathsf{r}}_{i},\hat{\mathsf{r}}^{c})\right|$. Hence, (27) and (28) are equivalent, that is, $\mathsf{z}$ in (27) is optimal if and only if $\mathsf{x}$ in (28) is optimal. ∎
@@
$w_{i}=\frac{\frac{1}{\sigma_{i}^{2}}}{\sum_{i=1}^{L}\frac{1}{\sigma_{i}^{2}}},\ i=1,2,\ldots,L.$ (32)
@@
$\hskip-10.00002pt\left\{\begin{aligned} &\mathrm{Re}(\hat{\mathsf{r}}^{c})=\arg\min\limits_{\mathrm{Re}(\mathsf{x})\in[0,M)}\sum_{i=1}^{L}w_{i}d_{M}^{2}\left(\mathrm{Re}(\tilde{\mathsf{r}}_{i}^{c}),\mathrm{Re}(\mathsf{x})\right),\\ &\mathrm{Im}(\hat{\mathsf{r}}^{c})=\arg\min\limits_{\mathrm{Im}(\mathsf{x})\in[0,M)}\sum_{i=1}^{L}w_{i}d_{M}^{2}\left(\mathrm{Im}(\tilde{\mathsf{r}}_{i}^{c}),\mathrm{Im}(\mathsf{x})\right).\end{aligned}\right$. (34)
@@
[…]
@@
$\Omega=\left\{\left\langle\sum_{i=1}^{L}w_{i}\tilde{\mathsf{r}}_{i}^{c}+M\left(\sum_{i=1}^{k_{1}}w_{\varsigma_{(i)}}+\mathrm{i}\sum_{i=1}^{k_{2}}w_{\upsilon_{(i)}}\right)\right\rangle_{M}:k_{1},k_{2}=1,2,\ldots,L\right\},$ (35)
where $\varsigma$ and $\upsilon$ are permutations on $\{1,2,\ldots,L\}$ satisfying
$\mathrm{Re}(\tilde{\mathsf{r}}_{\varsigma_{(1)}}^{c})\leq\mathrm{Re}(\tilde{\mathsf{r}}_{\varsigma_{(2)}}^{c})\leq\cdots\leq\mathrm{Re}(\tilde{\mathsf{r}}_{\varsigma_{(L)}}^{c})$
and
$\mathrm{Im}(\tilde{\mathsf{r}}_{\upsilon_{(1)}}^{c})\leq\mathrm{Im}(\tilde{\mathsf{r}}_{\upsilon_{(2)}}^{c})\leq\cdots\leq\mathrm{Im}(\tilde{\mathsf{r}}_{\upsilon_{(L)}}^{c}),$
respectively.
@@ IV-A Condition of Robust Estimation
We first consider a necessary condition of robust estimation for the MLE C-CRT. For convenience, we define the remainder error set as
$\mathcal{U}=\left\{\Delta\mathsf{r}_{1},\Delta\mathsf{r}_{2},\ldots,\Delta\mathsf{r}_{L}\right\}$
and the weighted average of the remainder errors as
$\overline{\Delta\mathsf{r}}=\sum_{i=1}^{L}w_{i}\Delta\mathsf{r}_{i},$
where the weights $w_{i}$ are defined in (32). In [10], a necessary condition for a robust estimation of real numbers is
$\frac{M}{2}\leq\Delta r_{i}-(\hat{N}-N)<\frac{M}{2}.$
For complex numbers, we have the following necessary condition:
@@
$\hskip-9.49997pt\mathrm{Re}(\Delta\mathsf{r}^{c})=\mathrm{Re}(\overline{\Delta\mathsf{r}})+\begin{cases}M,&\text{if}~\mathrm{Re}(\mathsf{r}^{c}+\overline{\Delta\mathsf{r}})<0,\\ 0,&\text{if}~0\leq\mathrm{Re}(\mathsf{r}^{c}+\overline{\Delta\mathsf{r}})<M,\\ -M,&\text{if}~\mathrm{Re}(\mathsf{r}^{c}+\overline{\Delta\mathsf{r}})\geq M\end{cases}$ (41)
@@
$\hskip-9.79996pt\mathrm{Im}(\Delta\mathsf{r}^{c})=\mathrm{Im}(\overline{\Delta\mathsf{r}})+\begin{cases}M,&\text{if}~\mathrm{Im}(\mathsf{r}^{c}+\overline{\Delta\mathsf{r}})<0,\\ 0,&\text{if}~0\leq\mathrm{Im}(\mathsf{r}^{c}+\overline{\Delta\mathsf{r}})<M,\\ -M,&\text{if}~\mathrm{Im}(\mathsf{r}^{c}+\overline{\Delta\mathsf{r}})\geq M.\end{cases}$ (42)
@@
For $\mathsf{N}$ and $\mathsf{N}_{0}$ in (15), $\mathsf{N}_{0}\in\mathcal{H}$ if and only if $\mathsf{N}\in M\mathcal{H}=\{M\mathsf{h}:\mathsf{h}\in\mathcal{H}\}$, where
$\mathcal{H}=\{h_{1}+h_{2}\mathrm{i}:1\leq h_{1},h_{2}<\Gamma-1\}.$
@@
Arbitrarily choose a $\Delta\mathsf{r}_{i}\in\mathcal{U}$, and let $\mathcal{V}=\{\Delta\mathsf{r}_{i}\}$. Then, we obtain from (39) that
$\Delta\mathsf{r}_{i}-\sum_{j\neq i}\frac{w_{j}}{\sum_{j\neq i}w_{j}}\Delta\mathsf{r}_{j}\in\mathcal{S}_{M}.$
Since $\sum_{j\neq i}w_{j}=1-w_{i}$, we have
$\frac{1}{1-w_{i}}\Delta\mathsf{r}_{i}-\frac{1}{1-w_{i}}\overline{\Delta\mathsf{r}}\in\mathcal{S}_{M}.$
Hence,
$\Delta\mathsf{r}_{i}-\overline{\Delta\mathsf{r}}\in(1-w_{i})\mathcal{S}_{M}.$
Consequently,
$\left[\frac{\Delta\mathsf{r}_{i}-\overline{\Delta\mathsf{r}}}{M}\right]=0.$
It follows from (43) that $\hat{\mathsf{q}}_{i}=\mathsf{q}_{i}-k_{1}-k_{2}\mathrm{i}$. Based on the arbitrariness of $\Delta\mathsf{r}_{i}$, we can obtain from (30) that
@@
Note that $\bar{\mathsf{\gamma}}_{i}\mathsf{\gamma}_{i}\equiv 1\mod\mathsf{\Gamma}_{i}$ and $\bar{\mathsf{\gamma}}_{j}\mathsf{\gamma}_{j}\equiv 0\mod\mathsf{\Gamma}_{i}$ for $j\neq i$. Hence,
$\sum_{i=1}^{L}\bar{\mathsf{\gamma}}_{i}\mathsf{\gamma}_{i}\equiv 1\mod\mathsf{\Gamma}_{i}.$
Consequently,
$\sum_{i=1}^{L}\bar{\mathsf{\gamma}}_{i}\mathsf{\gamma}_{i}(k_{1}+k_{2}\mathrm{i})\equiv k_{1}+k_{2}\mathrm{i}\mod\mathsf{\Gamma}_{i},$
that is, $\mathsf{\Gamma}_{i}$ divides $\sum_{i=1}^{L}\bar{\mathsf{\gamma}}_{i}\mathsf{\gamma}_{i}(k_{1}+k_{2}\mathrm{i})-(k_{1}+k_{2}\mathrm{i})$. Since $\mathsf{\Gamma}_{1},\mathsf{\Gamma}_{2},\ldots,\mathsf{\Gamma}_{L}$ are pairwise coprime, $\prod_{i=1}^{L}\mathsf{\Gamma}_{i}$ divides $\sum_{i=1}^{L}\bar{\mathsf{\gamma}}_{i}\mathsf{\gamma}_{i}(k_{1}+k_{2}\mathrm{i})-(k_{1}+k_{2}\mathrm{i})$. Thus,
$\sum_{i=1}^{L}\bar{\mathsf{\gamma}}_{i}\mathsf{\gamma}_{i}(k_{1}+k_{2}\mathrm{i})\equiv k_{1}+k_{2}\mathrm{i}\mod\Gamma.$
By (44), we have
$\hat{\mathsf{N}}_{0}\equiv\mathsf{N}_{0}-k_{1}-k_{2}\mathrm{i}\mod\Gamma.$
Since $\hat{\mathsf{N}}_{0}\in\mathcal{F}_{\Gamma}$, we have
$\hat{\mathsf{N}}_{0}=\left\langle\mathsf{N}_{0}-k_{1}-k_{2}\mathrm{i}\right\rangle_{\Gamma}.$
If $\mathsf{N}\in M\mathcal{H}$, i.e., $\mathsf{N}_{0}\in\mathcal{H}$ by Proposition 3, then $\hat{\mathsf{N}}_{0}=\mathsf{N}_{0}-k_{1}-k_{2}\mathrm{i}$. Consequently,
$\hat{\mathsf{N}}=M(\mathsf{N}_{0}-k_{1}-k_{2}\mathrm{i})+\mathsf{r}^{c}+\Delta\mathsf{r}^{c}=\mathsf{N}+\overline{\Delta\mathsf{r}}.$
Therefore, $\mathsf{N}$ can be robustly estimated. The next theorem demonstrates that (39) is both a necessary and sufficient condition for robust estimation.
@@
Theorem 6 demonstrates that the MLE C-CRT is capable of “preserving errors”, i.e., it preserves the weighted average error $\overline{\Delta\mathsf{r}}$ of the original remainder errors $\Delta\mathsf{r}_{1},\Delta\mathsf{r}_{2},\ldots,\Delta\mathsf{r}_{L}$ during the reconstruction. In this case, it is called error preserving MLE C-CRT. Since the output error of the MLE C-CRT is comparable to the input remainder error level, it is robust. By Theorem 6, we can derive the following robustness as well.
@@
Since $\mathrm{Re}(\Delta\mathsf{r}_{i})<\tau$ and $\mathrm{Im}(\Delta\mathsf{r}_{i})<\tau$, we have
$\left|\sum_{\Delta r_{i}\in S}\frac{w_{i}\mathrm{Re}(\Delta\mathsf{r}_{i})}{\sum_{\Delta r_{j}\in S}w_{j}}-\sum_{\Delta r_{i}\in\overline{S}}\frac{w_{i}\mathrm{Re}(\Delta\mathsf{r}_{i})}{\sum_{\Delta r_{j}\in\overline{S}}w_{j}}\right|<\left|\sum_{\Delta r_{i}\in S}\frac{w_{i}\tau}{\sum_{\Delta r_{j}\in S}w_{j}}\right|+\left|\sum_{\Delta r_{i}\in\overline{S}}\frac{w_{i}\tau}{\sum_{\Delta r_{j}\in\overline{S}}w_{j}}\right|=2\tau.$
Similarly,
$\left|\sum_{\Delta r_{i}\in S}\frac{w_{i}\mathrm{Im}(\Delta\mathsf{r}_{i})}{\sum_{\Delta r_{j}\in S}w_{j}}-\sum_{\Delta r_{i}\in\overline{S}}\frac{w_{i}\mathrm{Im}(\Delta\mathsf{r}_{i})}{\sum_{\Delta r_{j}\in\overline{S}}w_{j}}\right|<2\tau.$
By Theorem 6, we have $|\mathrm{Re}(\hat{\mathsf{N}})-\mathrm{Re}(\mathsf{N})|=|\mathrm{Re}(\overline{\Delta\mathsf{r}})|<\tau$ and $|\mathrm{Im}(\hat{\mathsf{N}})-\mathrm{Im}(\mathsf{N})|=|\mathrm{Im}(\overline{\Delta\mathsf{r}})|<\tau$. ∎